OpenStudy (anonymous):

How you write an equation of a line with a slope of -5 and a y-intercept of (0,3)? Do you use Slope-Intercept form?

OpenStudy (amistre64):

slpe intercept will work in this case yes

OpenStudy (anonymous):

You. Are. A. Saint. Thank you So much.

OpenStudy (anonymous):

let me know if the video helped any

OpenStudy (anonymous):

@amistre64 For two points, how would i go about putting it in an equation?

OpenStudy (anonymous):

and it helped Caleb. :)

OpenStudy (anonymous):

well it was supposed to be a video http://dradqrf3hyj25.cloudfront.net/thumbnails/Alg1_10_1_10-diagram_thumb.png

OpenStudy (anonymous):

Its more of an image. :P

OpenStudy (anonymous):

i am glad it did trying ti get the video version

OpenStudy (amistre64):

well, if we use the slope intercept form .. y = mx+b its just a matter of filling in the parts

OpenStudy (anonymous):

@amistre64 so say for instance you have (4,8) and (-2,6). how would you plug that in?

OpenStudy (amistre64):

well, when we dont have a slope given to us, we determine the slope using what we like to call: the slope formula the slope formula has a nice touch to it in that it can be reworked into the point slope eqaution of a line, which can then be worked into a slope intercept . given 2 points (x1,y1) and (x2,y2) y1 - y2 slope = ------ x1 - x2 now, knowing that the slope of a line is the same between any 2 points on the line, we can make one of the points general, just some (x,y) therefore y - yo slope = ------ x - xo mutliply to rid the bottom y - yo = slope(x - xo) , but slope is to wordy and we just name it a m y - yo = m(x - xo) this is hte point slope form of a line now, we can take this and solve for y to convert it into slope intercept form y = m(x-xo) + yo y = mx -mxo + yo , such that yo - mxo is the y intercept, which we rename as y = mx + b

OpenStudy (anonymous):

Oooh ok. I get it, thank you so much. :)